-29 Y-D X-1 The electric field may be zero at approximately O X=-0.4 OX= - 1.0 O x = 1.0 OX= 0.33
Answer #1

-28 E2 +01 E > ok b BA 2=1 x=0 Point B between two where electric field in Dro. first we consider Charges Electre filld filleElectric field at x=1, r² ra 1 unit E KO KO po option is incorrect Now consider to left side a pointe charge of b. distance tElectric field at e due to charge - 20 (k) (20) Ac² с AC- Ao toc th Ez = 2kg (1+b)? Net field at C, E = E2 - Ei 2KO KO (1+6)